# NdLinear: Don’t Flatten! Building Superior Neural Architectures by Preserving N-D Structure

### v2

@@ 4 Experimental Results
To empirically validate NdLinear, we conducted comprehensive experiments assessing its parameter efficiency, performance across diverse applications, and scalability. This section details the experimental setups, key results, and discusses the derived insights, organized by application domain.
4.1 Experimental Overview
Our extensive empirical investigation assesses NdLinear’s impact on parameter efficiency, its operational versatility, and its scalability. We benchmark NdLinear by integrating it as a drop-in replacement for standard Linear layers within influential neural architectures across key domains, demonstrating its benefits through a rigorous and wide-ranging suite of tests:
  • Natural Language Processing (Section 4.2): Our NLP evaluation is extensive, covering large language model pretraining (adapting OPT variants (Zhang et al., 2022)), text classification with BERT-based models (Devlin et al., 2019), and parameter-efficient finetuning. For the latter, we introduce and assess NdLinear-LoRA, showcasing its effectiveness on diverse mathematical reasoning and commonsense QA benchmarks with prominent models such as Qwen3-1.7B and LLaMA3-8B.
  • Time Series Analysis (Section 4.3): We investigate NdLinear’s performance in RNN-based models (Hochreiter and Schmidhuber, 1997) and Transformer-based models (Zerveas et al., 2021) for both forecasting and classification tasks. These evaluations span ten distinct datasets from the widely-used ETT and UCR archives.
  • Tabular Data (Section 4.4): We apply NdLinear in MLPs for tabular classification and regression tasks (Shwartz-Ziv and Armon, 2022), demonstrating its utility on this common data modality.
  • Computer Vision (Section 4.5): We evaluate NdLinear across a spectrum of vision tasks, including Vision Transformers (ViTs) (Dosovitskiy et al., 2021) for model distillation, CNNs (Krizhevsky et al., 2012b) for image classification (with detailed weight behavior analysis), and Diffusion Transformers (DiTs) (Peebles and Xie, 2023) for generative modeling.
This diverse set of experiments, encompassing varied architectures and numerous datasets per domain, allows for a thorough assessment of NdLinear. In all evaluations, NdLinear directly substitutes conventional Linear layers in critical network components (e.g., feature extractors, MLP blocks, timestep encoders, projection layers, and low-rank adapters). See more details and our fully reproducible code in the Appendix C.
4.2 Natural Language Processing Applications
4.2.1 Parameter-Efficient Finetuning with LoRA and NdLinear-LoRA
Low-rank adaptation (LoRA) (Hu et al., 2022) is a popular parameter-efficient strategy for finetuning LLMs. We investigate if a NdLinear layer, termed NdLinear-LoRA, can effectively replace the standard LoRA architecture (composed of two rank-$r$ matrices) while maintaining or improving downstream task performance with even fewer trainable parameters.
@@ 4.2.1 Parameter-Efficient Finetuning with LoRA and NdLinear-LoRA
Setup. We fine-tune two LLMs, Qwen3-1.7B (Yang et al., 2025) and LLaMA3-8B (Dubey et al., 2024). Training utilizes the OpenMathInstruct-1 (math10k) corpus (Toshniwal et al., 2024) for mathematical reasoning and the CommonsenseQA training split (Talmor et al., 2019) for commonsense reasoning.
Results. Table 1 presents accuracy on Math benchmarks (GSM8K (Cobbe et al., 2021), MultiArith (Roy and Roth, 2016)) and Commonsense Reasoning benchmarks (CSQA (Talmor et al., 2019), ARC-easy, ARC-challenge (Clark et al., 2018), BoolQ (Clark et al., 2019)). NdLinear-LoRA generally achieves comparable or superior accuracy to LoRA with significantly fewer trainable parameters (4x to 9x reduction compared to LoRA r=8).
  • Qwen3-1.7B: With only 1.15M trainable parameters (vs. 8.72M for LoRA r=8), NdLinear-LoRA shows substantial gains on GSM8K (+11.9 points over LoRA r=8) and improves performance on MultiArith. It achieves results comparable to LoRA r=8 on CSQA and ARC-e, and slightly trails on ARC-c.
  • LLaMA3-8B: Using only 2.26M parameters (vs. 20.97M for LoRA r=8, a 9$\approx$ 9$\times$ 9 × reduction), NdLinear-LoRA matches or exceeds LoRA r=8 on CSQA and ARC-c, and LoRA r=4 on ARC-e. While performance on GSM8K and MultiArith is lower than LoRA, the results are achieved with considerably fewer parameters.
Discussion.
These experiments, utilizing a straightforward implementation of NdLinear-LoRA, suggest that NdLinear’s structured transformations offer an effective alternative to LoRA’s low-rank matrix factorization. This presents a promising avenue for developing even more memory-efficient adaptation techniques for LLMs without performance degradation.
@@ Discussion.
Table 1: NdLinear-LoRA demonstrates significant parameter efficiency for LLM finetuning, achieving comparable or improved accuracy over standard LoRA methods with up to 9×9$\times$9 × fewer trainable parameters. Accuracy ($\uparrow$) after LoRA of Qwen3-1.7B and LLaMA3-8B models.
@@
Model Method Params GSM8K M.Arith CSQA ARC-e ARC-c BoolQ
Qwen3-1.7B LoRA ($r{=}4$) 4.36M 45.6 88.9 80.4 91.9 79.4 79.7
LoRA ($r{=}8$) 8.72M 40.3 82.2 80.9 91.8 79.3 80.8
NdLinear-LoRA 1.15M 52.2 90.0 81.0 92.2 78.3 79.7
LLaMA3-8B LoRA ($r{=}4$) 10.48M 50.5 84.4 80.6 90.4 76.3 85.1
LoRA ($r{=}8$) 20.97M 51.6 81.1 81.7 89.0 73.6 76.5
NdLinear-LoRA 2.26M 40.2 80.0 82.9 90.9 76.6 80.5
4.2.2 Efficient Large Language Model Pretraining
To evaluate NdLinear’s effectiveness at scale, we adapted the Open Pretrained Transformer (OPT) architecture (Zhang et al., 2022) by replacing Linear layers with NdLinear in the feedforward blocks.
Setup. We evaluated two variants: OPT-Small (124M parameters) and OPT-Mid (350M parameters). For these proof-of-concept experiments, both were trained from scratch using a subset of the original training corpus (BookCorpus and Wiki40B-English).
Results. Despite having fewer parameters, both OPT models with NdLinear layers achieved lower perplexity (Table 4 in Appendix D). The improvement was more pronounced in the larger model (perplexity gap increased from 0.215 in OPT-Small to 0.361 in OPT-Mid), suggesting that NdLinear’s benefits may scale with model size. The training and evaluation loss curves (Figure 1) further validate these improvements, showing consistently lower loss for NdLinear variants.
4.2.3 Efficient Text Classification with BERT
We evaluated NdLinear for text classification using BERT (Devlin et al., 2019) on the SST-2 (Socher et al., 2013) and CoLA (Warstadt et al., 2018) datasets.
@@ 4.2.3 Efficient Text Classification with BERT
Setup. In the BERT model, we replaced the conventional two-layer Linear classification head with a single NdLinear layer followed by a standard Linear classification layer.
@@
Results. As shown in Table 6 in Appendix D , NdLinear improved both accuracy and ROC AUC on both datasets while reducing the parameter count in the classification head by approximately 85%.
4.3 Time Series Analysis Applications
We evaluated NdLinear’s effectiveness in time series analysis, focusing on multivariate forecasting with RNNs and Transformers, and on classification tasks with Transformer-based models.
4.3.1 Efficient Multivariate Time Series Forecasting
We tested NdLinear in RNNs and Transformers for forecasting on four ETT (Electricity Transformer Temperature) datasets (Zhou et al., 2021).
Setup. The task was to predict the next 12 hours of data based on 24 hours of historical input. For RNNs, we compared a standard RNN against an RNN with NdLinear. For Transformers, we compared a 1-layer Transformer against a Transformer where Linear layers in the FeedForward blocks were replaced by NdLinear layers.
@@ 4.3.1 Efficient Multivariate Time Series Forecasting
Results. NdLinear consistently improved forecasting accuracy (lower MSE and MAE) across all datasets for both RNNs and Transformers (Table 3), while using significantly fewer parameters. The improvements were particularly notable on the ETTh1 and ETTh2 datasets.
4.3.2 Efficient Time Series Classification
Transformer models, while successful for time series classification, often have high computational demands due to their large parameter counts. We investigated NdLinear’s potential to mitigate this by enhancing parameter efficiency while maintaining or improving performance.
Setup. We applied Transformer-based models to six time series classification datasets from the UCR Archive (Dau et al., 2019). We replaced the standard Linear layers in the Transformer blocks with NdLinear layers. For a fair comparison, we also included a smaller standard Linear variant (denoted Base Model) where its hidden dimension matches the first transformation dimension of the corresponding NdLinear layer. F1 scores are reported in Table 7 in Appendix D.
@@ 4.3.2 Efficient Time Series Classification
Table 3: NdLinear substantially improves time-series forecasting accuracy and dramatically reduces parameter counts in both RNN and Transformer models on ETT datasets. The performance comparison on four ETT (Electricity Transformer Temperature) datasets. With NdLinear, model parameters are reduced by 50%, alongside consistent gains in accuracy. Metrics reported are mean $\pm$ standard deviation over three experimental runs.
@@
Results. NdLinear reduces parameter counts by up to 47% compared to the Linear layers in Transformers, while consistently improving F1 scores across all six datasets. Notably, even the standard Linear model with nearly twice the parameters often lags behind the NdLinear variant, highlighting NdLinear’s superior parameter efficiency and representational capacity for these tasks. The Linear variant, despite having a similar number of parameters to NdLinear, generally underperforms NdLinear, further emphasizing the architectural benefits of the NdLinear approach.
4.4 Tabular Data Applications
We evaluated NdLinear’s performance with MLPs on tabular data for both classification and regression tasks, demonstrating its ability to improve efficiency and performance in this common data modality.
Setup. We designed an MLP comprising two feature-extraction layers followed by a final linear layer for prediction. Within the feature-extraction layers, we compared blocks using standard Linear layers against blocks using NdLinear layers. Experiments were conducted on the Cardio Disease dataset (Sulianova, 2025) for classification and the Food Delivery Time dataset (Kumar, 2025) for regression.
Results. As shown in Table 7 in Appendix D, NdLinear demonstrated superior performance on both datasets while significantly reducing the number of parameters. For the Cardio Disease classification task, NdLinear achieved higher accuracy (0.7321 vs. 0.7265) with 67% fewer parameters. In the Food Delivery Time regression task, NdLinear yielded a lower MSE (67.824 vs. 70.508) with 58% fewer parameters.
4.5 Computer Vision Applications
We evaluated NdLinear in several computer vision tasks, including image classification with both ViTs and CNNs, and generative tasks with DiTs, to demonstrate its parameter efficiency and performance.
4.5.1 Parameter-Efficient Vision Transformer Distillation
@@
Setup. We explore NdLinear in ViTs (Dosovitskiy et al., 2021) for model distillation. A pre-trained ViT-B/16 model from (PyTorch, ) serves as the teacher. Student models, termed NdViT, are created by replacing all Linear layers in the feedforward blocks of a standard ViT with NdLinear layers. We experiment with various configurations, altering student model dimensions and the number of transformer blocks, evaluating on CIFAR-10 and CIFAR-100.
@@
Results. NdViT consistently outperforms conventional ViT baselines while using 26-68% fewer parameters (Table 9 and Figure 2 in Appendix D). Notably, an NdViT with embedding dimension 200 surpasses a naive ViT with dimension 500, highlighting NdLinear’s parameter efficiency. The performance advantage of NdViT tends to increase with more transformer blocks and larger dimensions.
4.5.2 Efficient Image Classification with CNNs
Setup. We trained a CNN on CIFAR-10 and CIFAR-100 datasets (Krizhevsky et al., 2009). The architecture consists of two convolutional layers followed by either a standard Linear layer or an NdLinear layer, and one final Linear layer as the classification head.
@@ 4.5.2 Efficient Image Classification with CNNs
Results. NdLinear achieves higher accuracy with significantly fewer parameters on both datasets (Table 8 in Appendix D). On CIFAR-10, NdLinear improves accuracy by 2.6% while reducing parameters by 94%. On CIFAR-100, it improves accuracy by 5.1% with 60% fewer parameters.
4.5.3 Efficient Generative Modeling with Diffusion Transformers (DiT)
We evaluated NdLinear in DiTs by training models from scratch on ImageNet-100 (Russakovsky et al., 2015) and by modifying pre-trained DiT models from (Jin and Xie, 2024). For image generation, FID (Dhariwal and Nichol, 2021) was the evaluation metric.
Training DiT from Scratch
To assess NdLinear’s effectiveness when training from scratch, we aimed to achieve superior performance with a comparable number of parameters. We replaced the two Linear layers in the timestep embedder of the standard DiT with NdLinear layers.
@@ 4.5.3 Efficient Generative Modeling with Diffusion Transformers (DiT)
Results. NdLinear variants consistently achieve lower (better) FID-10k scores compared to baseline DiTs when trained from scratch with similar parameter counts (Figure 3 in Appendix D).
@@
Results. Table 10 in Appendix D shows that NdLinear-based models (619M and 563M) achieve FID scores comparable to the larger baseline 674M model, despite significant parameter reductions.

### v3

@@ 4 Experimental Results
We evaluate NdLinear as a drop-in replacement for vanilla linear layers across language, vision, time-series, and tabular domains. Our experiments span CNNs, RNNs, Transformers (including LLMs), MLPs, and DiTs, with model scales from 65K to 8B parameters on over 20 datasets. We also compare NdLinear against alternative structured layers including Tensor Regression Layers (TRL/TCL) and Tensor-Train (TT) decompositions (Novikov et al., 2015; Kossaifi et al., 2020)
Universal findings: NdLinear reduces parameters by 50-95% while maintaining or improving performance across all tested configurations. When compared to other structured approaches, NdLinear consistently achieves superior accuracy with lower computational overhead. We present domain-specific results below, with ablations in Section  4.5 and full details in Appendix E.
4.1 Natural Language Processing
4.1.1 Parameter-Efficient Finetuning with NdLinear-LoRA
We introduce NdLinear-LoRA, replacing LoRA’s low-rank matrices $A,B$ with a single NdLinear adapter. Concretely, the LoRA update $\Delta W=BA$ is swapped for an NdLinear module that applies sequential mode-wise transforms on the reshaped activation tensor; all other LoRA mechanics (zero-init, scaling $\alpha$, residual addition, and merge-at-inference) remain unchanged. This preserves the native $N$-D structure during adaptation while keeping the same drop-in interface and training pipeline as standard LoRA.
@@ 4.1.1 Parameter-Efficient Finetuning with NdLinear-LoRA
We fine-tuned Qwen3-1.7B (Yang et al., 2025) and LLaMA3-8B (Dubey et al., 2024) on mathematical reasoning (OpenMathInstruct-1 (Toshniwal et al., 2024)) and commonsense QA (CommonsenseQA training split (Talmor et al., 2019)), then evaluated on GSM8K, MultiArith, CSQA, ARC-Easy, ARC-Challenge, and BoolQ..
Results. Table 1 shows NdLinear-LoRA achieves superior performance with dramatically fewer parameters. On Qwen3-1.7B with 87% fewer parameters (1.15M vs 8.72M), it improves GSM8K by 11.9 points and MultiArith by 7.8 points over standard LoRA. On LLaMA3-8B with 9× fewer parameters (2.26M vs 20.97M), it achieves the best CSQA (82.9%) and ARC-Challenge (76.6%) scores. This suggests structured transformations capture task-specific patterns more efficiently than low-rank factorization.
@@
Table 1: NdLinear-LoRA demonstrates significant parameter efficiency for LLM finetuning, achieving comparable or improved accuracy over standard LoRA methods with up to $9\times$ fewer trainable parameters. Accuracy ($\uparrow$) after LoRA of Qwen3-1.7B and LLaMA3-8B models.
@@
Model Method Params GSM8K M.Arith CSQA ARC‑e ARC‑c BoolQ
Qwen3‑1.7B LoRA (r=4r{=}4) 4.36M 45.6 88.9 80.4 91.9 79.4 79.7
LoRA (r=8r{=}8) 8.72M 40.3 82.2 80.9 91.8 79.3 80.8
NdLinear‑LoRA 1.15M 52.2 90.0 81.0 92.2 78.3 79.7
LLaMA3‑8B LoRA (r=4r{=}4) 10.48M 50.5 84.4 80.6 90.4 76.3 85.1
LoRA (r=8r{=}8) 20.97M 51.6 81.1 81.7 89.0 73.6 76.5
NdLinear‑LoRA 2.26M 40.2 80.0 82.9 90.9 76.6 80.5
4.1.2 Language Model Pretraining
We pretrained OPT models (Zhang et al., 2022) (124M and 350M parameters) from scratch on BookCorpus and Wiki40B-English, replacing feedforward linear layers with NdLinear. Despite fewer parameters, NdLinear variants achieve lower perplexity (Table 5 in the Appendix). The improvement scales with model size: perplexity gap increases from 0.215 (OPT-Small) to 0.361 (OPT-Mid), suggesting NdLinear’s benefits grow with scale.
4.1.3 Text Classification
@@
We evaluated NdLinear in BERT’s classification head (Devlin et al., 2019) on SST-2 sentiment analysis (Socher et al., 2013) and CoLA grammatical acceptability (Warstadt et al., 2018) tasks. We replaced BERT’s standard two-layer classification head with a single NdLinear layer followed by a final linear projection.
@@
Results. As shown in Table 7 in Appendix E, NdLinear improves both accuracy and ROC-AUC on both datasets while reducing the classification head parameters by approximately 85%. On SST-2, accuracy improves from 88.72% to 89.33%; on CoLA, from 77.90% to 79.06%.
4.2 Time Series Analysis
Multivariate Time Series Forecasting. We evaluated NdLinear in RNNs and Transformers for 12-hour ahead forecasting on four ETT (Electricity Transformer Temperature) datasets (Zhou et al., 2021), using 24 hours of historical input. For RNNs, we replaced vanilla linear layers with NdLinear. For Transformers, we replaced linear layers in feedforward blocks with NdLinear.
@@ 4.2 Time Series Analysis
Results. Table 3 shows NdLinear consistently improves forecasting accuracy across all datasets while using 50% fewer parameters. For RNNs on ETTh1, MSE decreases from 0.290 to 0.088 (70% reduction). For Transformers on ETTh2, MSE decreases from 0.0226 to 0.0158 (30% reduction). These improvements demonstrate NdLinear’s natural alignment with temporal-feature decomposition in multivariate time series.
Time Series Classification. We applied Transformer models with NdLinear to six UCR time series classification datasets (Dau et al., 2019), replacing vanilla linear layers in transformer blocks with NdLinear layers.
@@
Table 3: NdLinear substantially improves time-series forecasting accuracy and dramatically reduces parameter counts in both RNN and Transformer models on Electricity Transformer Temperature (ETT) datasets. With NdLinear, model parameters are reduced by 50%, alongside consistent gains in accuracy. Metrics reported are mean $\pm$ standard deviation over three experimental runs.
@@
Results. Table 8 in Appendix E shows NdLinear reduces parameters by up to 47% while consistently improving F1 scores across all datasets. For example, on LSST, F1 improves from 0.5928 to 0.6486 with 47% fewer parameters. Even compared to parameter-matched baselines (Base Model*), NdLinear variants achieve superior performance, highlighting that the architectural advantage extends beyond mere parameter reduction.
4.3 Tabular Data
We evaluated NdLinear in MLPs on tabular classification (Cardio Disease dataset) (Sulianova, 2025) and regression (Food Delivery Time dataset) (Kumar, 2025). We replaced vanilla linear layers in two-layer feature extractors with NdLinear layers.
Results. Table 2 in shows NdLinear improves performance while dramatically reducing parameters. For Cardio Disease classification, NdLinear achieves higher accuracy (73.21% vs 72.65%) with 67% fewer parameters. For Food Delivery Time regression, NdLinear reduces MSE from 70.51 to 67.82 with 58% fewer parameters. These results demonstrate that NdLinear’s dimensional factorization aligns well with tabular data where features often have independent effects.
4.4 Computer Vision
4.4.1 Vision Transformer Distillation
@@
We created student ViTs (NdViT) by replacing all linear layers in feedforward blocks with NdLinear, distilling from a pretrained ViT-B/16 teacher (Dosovitskiy et al., 2021). We tested various embedding dimensions (200, 300, 400) and depths (3, 6, 9 blocks) on CIFAR-10 and CIFAR-100 (Krizhevsky et al., 2009).
@@
Results. Table 10 and Figure 3 in Appendix E show NdViT consistently outperforms standard ViT students while using 26-68% fewer parameters. Notably, NdViT with embedding dimension 200 surpasses a standard ViT with dimension 500, demonstrating superior parameter efficiency. The performance gap increases with model depth, suggesting NdLinear’s benefits compound in deeper architectures.
4.4.2 CNN Image Classification
We replaced the penultimate linear layer in CNNs with NdLinear on CIFAR-10 and CIFAR-100, and compared against Tensor Regression Layers (TRL/TCL) and Tensor-Train (TT) layers.
@@ 4.4.2 CNN Image Classification
Results. NdLinear achieves higher accuracy with significantly fewer parameters on both datasets (Table 9 in Appendix E). On CIFAR-10, NdLinear improves accuracy by 2.6% while reducing parameters by 94%. On CIFAR-100, it improves accuracy by 5.1% with 60% fewer parameters.
Comparison with Structured Methods. Table 4 compares NdLinear against TRL/TCL and TT layers on CIFAR-100. NdLinear achieves superior accuracy (71.3%) compared to TRL/TCL (69.4%) and TT (56.2%), while using fewer parameters (434K vs 548K vs 769K), lower FLOPs (0.84G vs 3.97G vs 5.25G), 6× lower latency than TT (0.98ms vs 5.87ms), and minimal GPU memory overhead (35.16 MB vs 35.60 MB vs 100.44 MB). This demonstrates NdLinear’s practical advantages over alternative structured approaches.
Table 4: NdLinear outperforms existing structured tensor methods. Comparison with Tensor Regression Layers (TRL/TCL) and Tensor-Train decomposition on CIFAR-100 classification, replacing the CNN’s penultimate linear layer.
Method Mem (MB) Acc@5 Latency (s) FLOPs (G) Params
NdLinear 35.16 0.7133 0.000976 0.843 433,588
TRL/TCL 35.60 0.6935 0.001116 3.97 548,032
$T$ 100.44 0.5617 0.005871 5.25 769,316
4.4.3 Generative Modeling with Diffusion Transformers (DiT)
We evaluated NdLinear in DiTs through two approaches: training from scratch on ImageNet-100 (Russakovsky et al., 2015) and modifying pre-trained DiT models from (Jin and Xie, 2024). For training from scratch, we replaced linear layers in the timestep embedder. For pretrained models, we replaced linear layers in both timestep embedders and attention MLPs.
@@ 4.4.3 Generative Modeling with Diffusion Transformers (DiT)
Results. When trained from scratch, NdLinear variants achieve lower (better) FID scores with comparable parameters (Figure 4 in Appendix E). For pretrained models, NdLinear reduces parameters from 674M to 619M while maintaining similar FID scores. demonstrating that parameter reduction doesn’t sacrifice generation quality.
@@
Results. Table 11 in Appendix E shows that NdLinear-based models (619M and 563M) achieve FID scores comparable to the larger baseline 674M model, despite significant parameter reductions.
4.5 Ablation Studies
We conducted extensive ablations to understand NdLinear’s design choices and practical considerations (full details in Appendix F).
4.5.1 Design Choices
Per-mode Bias Terms. Including bias terms for each dimension transformation significantly improves performance, with benefits increasing at larger widths. On the Radius Bump task, bias terms improve MSE by 4.5% at width 32 and 15.2% at width 128, with negligible parameter overhead (<5%) (Table F.1 in the Appendix). This suggests per-mode biases help capture dimension-specific offsets. More information on the taks can be found in Appendix F.3
Axis Ordering Robustness. NdLinear shows remarkable robustness to transformation order. Permuting axes (original, reverse, random) cause only 4% accuracy variation on CIFAR-100, with random ordering achieving 96% of baseline performance (Table F.2 in the Appendix). This suggests NdLinear learns relatively axis-independent features, simplifying deployment.
Hidden dimensions offer a direct trade-off between efficiency and expressivity. Doubling hidden size improved accuracy by 2% while still using 40% fewer parameters than dense layers.
4.5.2 Practical Considerations
Hyperparameter Robustness. NdLinear requires no special tuning. Across learning rates (0.001, 0.01), batch sizes (32, 64, 128), and hidden configurations, accuracy varies only by 11 percentage points (64-75%) on CIAFR-10 (Table F.4 in the Appendix). Even the worst configuration (LR=0.01, batch=32) achieves 64% accuracy with 6× fewer parameters than dense layers, demonstrating robustness to suboptimal settings.
Computational Overhead. Despite sequential processing, measured overheads are minimal: peak memory increases by 1.1-2.0% (CIFAR CNN: 35.17→36.91 MB), training time by 0.6-1.6% (47.2→47.8s per epoch). Inference latency remains comparable due to reduced FLOPs offsetting sequential operations (Table F.6 in the Appendix).
Sample Efficiency. The separability bias dramatically affects data efficiency (Section F.5). On our synthetic tasks with $\alpha=0.1$ separability, NdLinear reaches target error with 5× fewer samples (2,000 vs 10,000). Conversely, at $\alpha=0.9$ entanglement, it requires 1.7× more samples (25,000 vs 15,000), quantifying the bias-variance trade-off. The representation analysis in Section 3.1 provides additional mechanistic insight into these efficiency gains.
Architectural Comparisons. As shown in Section 4.4.2, NdLinear outperforms alternative structured layers (TRL/TCL, TT) on all metrics, validating our design choices.
Key Takeaway.
NdLinear is robust to implementation details and requires no special tuning, where standard hyperparameters work well. The critical decision is architectural: whether the task exhibits sufficient separability to benefit from NdLinear’s bias.